# In-Context Learning State Vector with Inner and Momentum Optimization

### v1

@@ Abstract
Large Language Models (LLMs) have exhibited an impressive ability to perform In-Context Learning (ICL) from only a few examples. Recent works have indicated that the functions learned by ICL can be represented through compressed vectors derived from the transformer. However, the working mechanisms and optimization of these vectors are yet to be thoroughly explored. In this paper, we address this gap by presenting a comprehensive analysis of these compressed vectors, drawing parallels to the parameters trained with gradient descent, and introduce the concept of state vector. Inspired by the works on model soup and momentum-based gradient descent, we propose inner and momentum optimization methods that are applied to refine the state vector progressively as test-time adaptation. Moreover, we simulate state vector aggregation in the multiple example setting, where demonstrations comprising numerous examples are usually too lengthy for regular ICL, and further propose a divide-and-conquer aggregation method to address this challenge. We conduct extensive experiments using Llama-2 and GPT-J in both zero-shot setting and few-shot setting. The experimental results show that our optimization method effectively enhances the state vector and achieves the state-of-the-art performance on diverse tasks. Code is available at https://github.com/HITsz-TMG/ICL-State-Vector
@@ 1 Introduction
In-Context Learning (ICL) has emerged as a powerful capability in tandem with the scaling of large language models (LLMs) (Brown et al., 2020). By instructing LLMs with few-shot demonstration examples, ICL enables these models to perform a diverse array of tasks and delivers remarkable empirical results in various downstream Natural Language Processing (NLP) tasks (Wei et al., 2022; Liu et al., 2023a). For example, a model prompted with the input “gaot goat, sakne snake, brid ” can produce the output “bird”. Given these successes, it is worthwhile to inquire about the exact internal working mechanisms of ICL. Considering the ambiguity surrounding the operation of ICL through the auto-regressive transformer, it might function as a general mechanism that utilizes both demonstrations and the query to yield the prediction (Dong et al., 2023).
@@
To gain a better understanding of ICL, previous works have studied its mechanisms from a theoretical viewpoint and offered several functional assumptions (Olsson et al., 2022; Xie et al., 2022; Garg et al., 2022; von Oswald et al., 2023; Dai et al., 2023). More recently, some studies have found that the ICL mapping function exists in the outputs of the attention layers or attention heads (Liu et al., 2023b; Dai et al., 2023) when applying causal effects analysis on a different set of models and tasks, such as the task vector (Hendel et al., 2023) and the function vector (Todd et al., 2023). These works show that the functionalities learned through ICL can be encapsulated in compressed vectors derived from transformers, which then can be used to intervene the transformer for handling queries without demonstrations. This revelation suggests a shift in understanding the learning dynamics of LLMs, where the focus is not just on the raw power of large-scale models but also on how efficiently they can be adapted. However, there remains a critical gap in understanding the operational mechanisms and optimization strategies of these compressed vectors. Unlike traditional parameters that undergo extensive training, these vectors represent a more nuanced aspect of ICL, one that encapsulates learned information in a more condensed form.
@@
In this paper, we aim to bridge the existing gap by presenting a comprehensive analysis of compressed vectors. Specifically, we investigate their similarities with parameters trained via gradient descent and introduce the state vector that encapsulates the processing state of ICL stored in the attention activation. To further understand how these vectors can be optimized and enhanced, drawing insights from the model soup (Wortsman et al., 2022) and momentum-based gradient optimization algorithm (Qian, 1999; Sutskever et al., 2013), we propose two methodologies (namely, inner optimization and momentum optimization) which are designed to incrementally enhance the state vector. Moreover, we delve into the practical challenges of applying ICL in the multiple example setting where demonstrations comprising numerous examples are typically too lengthy for regular ICL processes, such as in the 100-shot setting. To scale up for processing extended examples, we further introduce a divide-and-conquer aggregation method that effectively compresses the ICL function of these extensive examples into a state vector. We conduct extensive experiments using Llama-2 (Touvron et al., 2023) and GPT-J (Wang and Komatsuzaki, 2021) in both zero-shot and few-shot settings. The experimental results show that our method effectively enhances the state vector and we achieve state-of-the-art performance on diverse tasks. This not only indicates the effectiveness of our approach but also paves the way for more comprehensive understanding of ICL.
@@
We focus on understanding how compressed vectors of ICL work and provide a perspective that these vectors share similarities with parameters trained through gradient descent. Building on this observation, we then propose the concept of the state vector.
Inspired by the work on model soup and momentum-based gradient optimization algorithm, we propose inner and momentum optimization to refine the state vector as test-time adaptation. Moreover, we further introduce a divide-and-conquer approach to simulate state vector aggregation for scaling up to large numbers of examples.
We show that the proposed methods are empirically practical for a wide range of tasks through extensive experiments and provide implications for future research to fully characterize function execution by ICL.
@@ 2 Related Work
Recent works have focused on the working mechanisms of ICL (Chan et al., 2022; Xie et al., 2022; Wang et al., 2023). Olsson et al. (2022) argue that induction heads may be the mechanistic source of general ICL in transformers. Akyürek et al. (2022) show that transformer-based in-context learners can implicitly implement standard optimization algorithms on linear models. A mainstream assumption posits that ICL has similarity with the gradient descent. von Oswald et al. (2023) demonstrate how a linear attention-only transformer model can perform a gradient descent-like procedure implicitly. Dai et al. (2023) compare standard gradient descent based fine-tuning and ICL, and figure out that the transformer attention of ICL exhibits a dual form of gradient descent-based optimization. Moreover, some works revisit and modify this theory on the layer causality dependence (Natan et al., 2023) or training batch size (Shen et al., 2023). In contrast, we focus on the application of the dual form of gradient descent and ICL and present optimization methods with inspiration from the dual form.
@@
Numerous studies have explored the concept of compressing various tasks into task representations as a means of effectively manipulating tasks within ICL ability. Notably, Shao et al. (2023) and Mu et al. (2023) yield compositional task representations by training a composition model. In a slightly different vein, some researchers have delved into the art of devising methodologies to compose minor parameter adjustments acquired through task fine-tuning (Ilharco et al., 2022; Panigrahi et al., 2023; Yu et al., 2023; Hu et al., 2024; Merullo et al., 2023). An alternative line of research finds that the task representation could be extracted in ICL (Liu et al., 2023b; Hendel et al., 2023; Todd et al., 2023; Yang et al., 2023). Different from these approaches, our work avoids the need for additional training and focuses more on analysing why these compressed vectors work and how to improve their performance.
@@ 3 Formalization
Here we analysis the attention activation of the last separate token. In the -th transformer layer, the output activation of the attention heads of the last separate token is:
@@
With the above dual form between activation and trained parameters, in light of the observations that transformers tend to learn the ICL function in their first layers (Wang et al., 2023), we have the following hypothesis: During the process of ICL, the first layers progressively update the flow of information using each example in the demonstration through forward computation. The processing state of ICL is then stored within the activation of the attention head. The subsequent layers access the processing state to reinstate the ICL function, which is used implicitly for predicting the inputs. Therefore we concatenate the activation in the initial layers and introduce the notation of the state vector:
where is the number of layers and is the number of examples in the demonstration. denotes the concatenation operation. Note that we have a completely different construction strategy and usage compared to the function vector (Todd et al., 2023). Although the task vector (Hendel et al., 2023) may be functionally equivalent in the forward process, the proposed state vector differs significantly in its integration into the model, making it easier and more effective to analysis and interpret.
@@ 4.1 Overview
As illustrated in Figure 1, our approach initially extracts the state vector from the attention head that corresponds to the final separate token in the first layers using the demonstration and a dummy query. Then, considering the state vector as trained parameters, coupled with drawing inspiration from the model soup and the momentum-based gradient optimization algorithm, we introduce two methods that progressively optimize the state vector as test-time adaption (Liang et al., 2023): (1) inner optimization (§4.2) and (2) momentum optimization (§4.3). Moreover, we propose a divide-and-conquer (D&C) state vector aggregation for gathering the ICL function in the multiple example setting (§4.4).
After the state vector optimization or aggregation, we utilize the processed state vector to intervene the model during the forward inference pass. In particular, we first input a test query in the zero-shot setting or with the demonstration in the few-shot setting. During the forward pass in the first layers, we replace the attention activation of the final separate token with the corresponding activation in the state vector. In other words, the state vector is leveraged to intervene in the output of the first transformer layers, blocking the attention of the last separate token to the previous context. With state vector intervention, the transformer learns the ICL function from the processing state stored in the state vector, and continues to make the prediction on the test query.
@@ 4.2 Inner Optimization
Inspired by the works on the model soup (Wortsman et al., 2022; Chronopoulou et al., 2023) which show that weight-space averaging not only yields performance improvement but also often enhances robustness, we thus ask the following research question (RQ1): Is it possible to optimize our state vector using the model soup approach? To explore this question, we propose an inner optimization method to improve the effectiveness and robustness of state vector. Specifically, we not only extract the state vector in each separate token of the dummy query but also extract the state vector from each example. Within the derivation in §3, the state vector can be considered as parameters trained through the preceding example of its corresponding separate token. Formally, with a forward pass in an shot ICL setting, we can extract the state vector () from the +1-th separate token. Subsequently, we apply a uniform averaging process to these state vectors as follows:
@@ 4.3 Momentum Optimization
Since we view the state vector as parameters trained gradually through demonstration examples, the difference between two state vectors with adjacent corresponding separate tokens can also be regarded as the influence of the middle example, akin to the gradient. Motivated by this understanding, coupled with extensive studies of the gradient optimization algorithm (Sutskever et al., 2013; Duchi et al., 2010; Loshchilov and Hutter, 2019), we direct our focus toward a simple momentum-based gradient optimization algorithm, seeking to answer the following research question (RQ2): Can our state vector be optimized using momentum-based optimization algorithm? To answer this question, we propose a momentum optimization. Formally, we first extract the influence of each example by subtracting two adjacent state vectors.
where is the influence of -th () example in the early layer. Then, we apply the momentum gradient optimization algorithm to obtain optimized influence , and add it to the last state vector.
where is the momentum optimized state vector and is the inner optimized state vector. denotes the momentum gradient optimization algorithm. We also explore various other optimization algorithms in §7.1.
@@ 4.4 Divide-and-Conquer Aggregation
In addition to optimizing the state vector to more effectively represent the ICL function from a small number of examples, we also explore its capacity to encapsulate multiple examples within a single vector. However, regular ICL can not be directly used on multiple examples due to the context length limitation of current LLMs. This leads us to investigate the following question (RQ3): Can we use the state vector to represent multiple examples that are unmanageable for regular ICL? To address this question, we propose a divide-and-conquer method for state vector aggregation. As depicted in Figure 1, our approach involves distinct aggregation processes (i.e. the divide stage and the conquer stage). In the divide stage, we first divide the examples into groups, which we refer to as grouped demonstrations. Within each group, we insert a specific dummy query from which we extract a state vector, called the Group State Vector (GSV). Subsequently, in the conquer stage, we pair the dummy queries with their associated labels as an example to create an aggregated demonstration. Then by inserting a new dummy query to the aggregated demonstration, we are able to extract the aggregated state vector from it. This process includes a crucial step where in the forward pass, we replace the activations of each separate token in examples with their corresponding GSVs. This allows us to aggregate the ICL function of each grouped demonstration into its GSV, and subsequently aggregate the ICL function of each GSV into a single, comprehensive aggregated state vector. This aggregated vector is then utilized for interventions during inference, same as the usage of the optimized state vector in §4.2 and §4.3. Moreover, in the few-shot setting, the aggregated demonstrations are treated as few-shot demonstrations. The divide-and-conquer approach effectively circumvents the context-length constraints inherent in LLMs, thereby enabling more efficient aggregation of information across multiple examples.
5 Setup
@@
For each dataset, we use random sampling to create subsets. Each subset consists of 10 instances for demonstrations and one instance for a dummy query. The remaining instances are split into test and development sets with a 7:3 ratio. For experiment with multiple examples, we sample 100 instances instead of 10. For simplicity, we restrict to single-token output and use first output token accuracy as the evaluation metric as in previous work (Hendel et al., 2023; Todd et al., 2023).
@@
Regular is the baseline for the zero shot setting that uses only the given query as input, while ICL baseline (Wei et al., 2022) makes predictions on the label by taking both the demonstrations and the given query.
@@
Task vector (Hendel et al., 2023) is extracted from the hidden state of the separate token and are leveraged for blocking the layer when inference.
6 Results
@@ 6.1 Inner Optimization (RQ1)
As shown in Table 1, the performance of our inner optimized state vector has a significant improvement comparing the task vector and function vector in both zero-shot and few-shot settings. Our state vector with inner optimization. In the zero-shot setting, the inner optimization shows an average improvement of 10.2% on Llama-2 and 5.9% on GPT-J across six datasets. In the few-shot setting, the inner optimization also achieves 1.2% improvement on Llama-2 and 1.7% on GPT-J. The improvement demonstrates the effectiveness of inner optimization. However, although state vector (inn.) outperforms task vector, its few-shot performance on some datasets is inferior to the ICL baseline. We attribute this primarily to the introduction of query information from examples. While inner optimization enhances task-relevant information for the state vector, it also introduces noise of other dummy queries, hindering the model’s ability to focus on the current predictive query, thereby reducing performance. In addition to the performance improvements, our inner optimization approach also effectively alleviates the phenomenon of high variance in the original task vector in the zero-shot setting. In practical use, the performance of the task vector is influenced by demonstrations and dummy queries, leading to weaker robustness. Our proposed inner optimization approach effectively mitigates this issue, similarly motivated as the model averaging method, thereby enhancing the robustness of the state vector.

### v2

@@ Abstract
Large Language Models (LLMs) have exhibited an impressive ability to perform In-Context Learning (ICL) from only a few examples. Recent works have indicated that the functions learned by ICL can be represented through compressed vectors derived from the transformer. However, the working mechanisms and optimization of these vectors are yet to be thoroughly explored. In this paper, we address this gap by presenting a comprehensive analysis of these compressed vectors, drawing parallels to the parameters trained with gradient descent, and introducing the concept of state vector. Inspired by the works on model soup and momentum-based gradient descent, we propose inner and momentum optimization methods that are applied to refine the state vector progressively as test-time adaptation. Moreover, we simulate state vector aggregation in the multiple example setting, where demonstrations comprising numerous examples are usually too lengthy for regular ICL, and further propose a divide-and-conquer aggregation method to address this challenge. We conduct extensive experiments using Llama-2 and GPT-J in both zero-shot setting and few-shot setting. The experimental results show that our optimization method effectively enhances the state vector and achieves the state-of-the-art performance on diverse tasks. Code is available at https://github.com/HITsz-TMG/ICL-State-Vector
@@ 1 Introduction
In-Context Learning (ICL) has emerged as a powerful capability in tandem with the scaling of large language models (LLMs) (Brown et al., 2020). By simply conditioning on a few input-label pairs as demonstrations, LLMs yield a significant improvement and deliver remarkable results in various downstream Natural Language Processing (NLP) tasks (Wei et al., 2022; Liu et al., 2023a). For example, a model prompted with the input “gaot goat, sakne snake, brid ” can produce the output “bird”. Given these successes, it is worthwhile to inquire about the exact internal working mechanisms of ICL. Considering the opaque operation of ICL within the auto-regressive transformer, it is plausible that ICL might function as a general mechanism that leverages both demonstrations and the query to yield the prediction (Dong et al., 2023).
@@
Recently, some studies have found that the ICL mapping function exists in the outputs of the attention layers or attention heads (Liu et al., 2023b; Dai et al., 2023) when applying causal effects analysis on a different set of models and tasks, such as the task vector (Hendel et al., 2023) and the function vector (Todd et al., 2023). These works show that the functionalities learned through ICL can be encapsulated in compressed vectors derived from transformers, which then can be used to intervene in the transformer to handle queries without demonstrations. This revelation suggests the potential mechanism of ICL that first utilises demonstrations to learn the mapping function from inputs to labels in shallow transformer layers, and then uses the ICL function in deeper transformer layers to make predictions (Hendel et al., 2023). However, while these compressed vectors encapsulate learned information in a more condensed form and show significant promise in applying ICL, there still exists a considerable gap in understanding the operational mechanisms and optimization strategies of these vectors. This significant gap hinders the further grasping and utilization of ICL.
@@
In this paper, we aim to bridge the existing gap by presenting a comprehensive analysis of compressed vectors. Specifically, we investigate their similarities with parameters trained via gradient descent and introduce the formulation of state vector that encapsulates the processing state of ICL stored in the attention activations. Building on the concept of state vector, and drawing insights from the model soup (Wortsman et al., 2022) and momentum-based gradient optimization algorithms (Qian, 1999; Sutskever et al., 2013), we propose inner optimization and momentum optimization strategies which are progressively applied to enhance the state vector. Moreover, we further exploit the demonstration compression capabilities of the state vector to address the practical challenges encountered when applying ICL in settings with multiple examples, where demonstrations are typically too lengthy for standard ICL, such as in the 100-shot setting which is prevalent in practice. Specifically, we introduce a divide-and-conquer aggregation method that effectively aggregates the ICL functions of these extensive examples. This approach enables us to scale up for processing extended examples by compressing them into a single state vector. We conduct extensive experiments using Llama-2 (Touvron et al., 2023) and GPT-J (Wang and Komatsuzaki, 2021) in both zero-shot and few-shot settings. The experimental results show that our method effectively enhances the state vector and achieves state-of-the-art performance on diverse tasks. This not only manifests the effectiveness of our approach but also paves the way for a more comprehensive understanding of ICL.
@@
We delve into the working mechanism of compressed vectors in ICL and highlight their similarities with parameters trained via gradient descent. Building on this observation, we propose the formulation of the state vector.
We propose inner and momentum optimization to progressively refine the state vector as an efficient test-time adaptation. Additionally, we introduce a divide-and-conquer aggregation to effectively scale up to large numbers of examples.
We show the practicality of our proposed methods across a wide range of tasks through extensive experiments. Our results also offer insights for future research aiming to fully understand the functionalities of ICL.
@@ 2 Related Work
Recent works have focused on the working mechanisms of ICL (Chan et al., 2022; Xie et al., 2022; Wang et al., 2023). Olsson et al. (2022) argue that induction heads may be the mechanistic source of general ICL in transformers. Akyürek et al. (2022) show that transformer-based in-context learners can implicitly implement standard optimization algorithms on linear models. A mainstream assumption posits that ICL has a similarity with the gradient descent. von Oswald et al. (2023) demonstrate how a linear attention-only transformer model can perform a gradient descent-like procedure implicitly. Dai et al. (2023) compare standard gradient descent based fine-tuning and ICL, and figure out that the transformer attention of ICL exhibits a dual form of gradient descent-based optimization. Moreover, some works revisit and modify this theory on the layer causality dependence (Natan et al., 2023) or training batch size (Shen et al., 2023). In contrast, we focus on the application of the dual form of gradient descent and ICL and present optimization methods with inspiration from the dual form.
@@
Numerous studies have extensively explored the concept of compressing various tasks into task representations as a means of effectively manipulating tasks within ICL ability. Notably, Shao et al. (2023) and Mu et al. (2023) have successfully yielded compositional task representations by training a composition model. In a slightly different vein, some researchers have delved into the art of devising methodologies to compose minor parameter adjustments acquired through task fine-tuning (Ilharco et al., 2022; Panigrahi et al., 2023; Yu et al., 2023; Hu et al., 2024; Merullo et al., 2023). An alternative line of research finds that the task representation could be extracted in ICL (Liu et al., 2023b; Hendel et al., 2023; Todd et al., 2023; Yang et al., 2023). Different from these approaches, our work avoids the need for additional training and focuses more on analysing why these compressed vectors work and how to improve their performance.
@@ 3 Formalization
Here we analyse the attention activation of the last separate token. In the -th transformer layer, the output activation of the attention heads of the last separate token is:
@@
With the above dual form between activation and trained parameters, and in light of observations that transformers tend to learn the ICL function primarily in their first layers (Wang et al., 2023), we have the following hypothesis: During the process of ICL, the first layers progressively update the flow of information using each example in the demonstration through forward computation. The processing state of ICL is then stored within the activation of the attention head. The subsequent layers access and utilize the processing state to reinstate the ICL function, which is used implicitly for predicting the queries. Therefore we concatenate the activation in the initial layers and introduce the notation of the state vector:
where is the number of layers and is the number of examples in the demonstration. denotes the concatenation operation. Note that we have a completely different construction strategy and usage compared to the function vector (Todd et al., 2023). Although the task vector (Hendel et al., 2023) may be functionally equivalent in the forward process, the proposed state vector differs significantly in its integration into the model, making it easier and more effective to analyse and interpret.
@@ 4.1 Overview
As illustrated in Figure 1, our approach initially extracts the state vector from the attention head that corresponds to the final separate token in the first layers using a demonstration and a dummy query. Then, with the view of treating the state vector as trained parameters, coupled with drawing inspiration from the model soup and the momentum-based gradient optimization algorithm, we introduce two methods that progressively optimize the state vector as test-time adaptation (Liang et al., 2023): (1) inner optimization (§4.2) and (2) momentum optimization (§4.3). Moreover, we propose a divide-and-conquer (D&C) state vector aggregation method for efficiently compressing the ICL function in the multiple example setting (§4.4).
After the state vector optimization or aggregation, we utilize the processed state vector to intervene the model during the forward inference pass. In particular, we first input a test query in the zero-shot setting or with the demonstration in the few-shot setting. During the forward pass in the first layers, we replace the attention activation of the last separate token with the corresponding activation in the state vector. In other words, the state vector is leveraged to intervene in the output of the first transformer layers, blocking the attention of the last separate token to the previous context. With state vector intervention, the transformer learns the ICL function from the processing state stored in the state vector, and continues to make the prediction on the test query.
@@ 4.2 Inner Optimization
Inspired by the works on the model soup (Wortsman et al., 2022; Chronopoulou et al., 2023) which show that weight-space averaging not only yields performance improvement but also often enhances robustness, we thus ask the following research question (RQ1): Is it possible to optimize our state vector using the model soup approach? To explore this question, we propose an inner optimization method to improve the effectiveness and robustness of state vector. Specifically, we not only extract the state vector in each separate token of the dummy query but also extract the state vector from each example. Formally, with a forward pass in an shot ICL setting, we extract the state vector () from last separate token. Subsequently, we apply a uniform averaging process to these state vectors as follows:
@@ 4.3 Momentum Optimization
Since we view the state vector as parameters trained gradually through demonstration examples, the difference between two state vectors with adjacent corresponding separate tokens can also be regarded as the influence of the middle example, akin to the gradient. Motivated by this understanding, coupled with extensive studies of the gradient optimization algorithm (Sutskever et al., 2013; Duchi et al., 2010; Loshchilov and Hutter, 2019), we direct our focus toward a simple momentum-based gradient optimization algorithm, seeking to answer the following research question (RQ2): Can our state vector be optimized using momentum-based optimization algorithm? To answer this question, we propose a momentum optimization. Formally, we first extract the influence of each example by subtracting two adjacent state vectors:
where is the influence of -th () example in the early layer. Then, we apply the momentum gradient optimization algorithm to obtain optimized influence , and add it to the last state vector:
where is the momentum optimized state vector and is the inner optimized state vector. denotes the momentum gradient optimization algorithm. We also explore various other gradient optimization algorithms in §6.1.
@@ 4.4 Divide-and-Conquer Aggregation
In addition to optimizing the state vector to more effectively represent the ICL function from a small number of examples, we also explore its capacity to encapsulate multiple examples within a single vector. However, regular ICL can not be directly used on multiple examples due to the context length limitation of current LLMs. This leads us to investigate the following question (RQ3): Can we use the state vector to represent multiple examples that are unmanageable for regular ICL? To address this question, we propose a divide-and-conquer method for state vector aggregation. As depicted in Figure 1, our approach involves distinct aggregation processes (i.e. the divide stage and the conquer stage). In the divide stage, examples are randomly divided into groups, termed grouped demonstrations. Within each group, a random example is selected to serve as a dummy query, which allows us to extract a group-specific state vector. In the conquer stage, these dummy queries are paired with their corresponding labels to form input-label pairs. From these input-label pairs, we form an aggregated demonstration, add an additional dummy query, and subsequently extract the aggregated state vector. It is worth noting that during the forward pass of aggregated state vector extraction, we utilise the group-specific state vector to intervene the attention activation of the separate tokens of their corresponding examples. The divide and conquer approach allows us to aggregate the ICL function of each grouped demonstration into its respective group-specific state vector, and subsequently aggregate the ICL function of each group-specific state vector into a single, comprehensive aggregated state vector. This aggregated vector is then utilized for interventions during inference, similarly to the optimized state vector discussed in §4.2 and §4.3. Moreover, in the few-shot setting, the aggregated demonstrations are treated as inference demonstrations. The divide-and-conquer approach effectively circumvents the context-length constraints inherent in LLMs, thereby enabling a more effective and efficient aggregation of information across multiple examples.
5 Experiment
5.1 Setup
@@
We employ Llama-2-7B and GPT-J-6B as our LLMs, chosen for their moderate model sizes, open-source and capability for ICL. We also provide the results with larger models (i.e., Llama-2-13B) in the Appendix H. We use Llama-2-7B as the default model unless otherwise specified. Our method is orthogonal to the choice of transformer-based decoder-only autoregressive LLMs.
5.2 Baseline
@@
Regular is the baseline for the zero-shot setting that uses only the given query as input, while ICL baseline (Wei et al., 2022) makes predictions on the label by taking both the demonstrations and the given query.
@@
Task vector (Hendel et al., 2023) is extracted from the hidden state of the separate token and is leveraged for blocking the layer when inference.
@@ 5.3 Inner Optimization(RQ1)
As shown in Table 1, the performance of our inner optimized state vector has a significant improvement comparing the task vector and function vector in both zero-shot and few-shot settings. Our state vector with inner optimization. In the zero-shot setting, the inner optimization shows an average improvement of 10.2% on Llama-2 and 5.9% on GPT-J across six datasets. In the few-shot setting, the inner optimization also achieves a 1.2% improvement on Llama-2 and 1.7% on GPT-J. The improvement demonstrates the effectiveness of inner optimization. However, although state vector (inn.) outperforms task vector, its few-shot performance on some datasets is inferior to the ICL baseline. We attribute this primarily to the introduction of query information from examples. While inner optimization enhances task-relevant information for the state vector, it also introduces noise of other dummy queries, hindering the model’s ability to focus on the current predictive query, thereby reducing performance. In addition to the performance improvements, our inner optimization approach also effectively alleviates the phenomenon of high variance in the original task vector in the zero-shot setting. In practical use, the performance of the task vector is influenced by demonstrations and dummy queries, leading to weaker robustness. Our proposed inner optimization approach effectively mitigates this issue, similarly motivated as the model averaging method, thereby enhancing the robustness of the state vector.